# Improved Zero-Shot Classification by Adapting VLMs with Text Descriptions

### v1

@@ Abstract
The zero-shot performance of existing vision-language models (VLMs) such as CLIP [25] is limited by the availability of large-scale, aligned image and text datasets in specific domains. In this work, we leverage two complementary sources of information—descriptions of categories generated by large language models (LLMs) and abundant, fine-grained image classification datasets—to improve the zero-shot classification performance of VLMs across fine-grained domains. On the technical side, we develop methods to train VLMs with this “bag-level” image-text supervision. We find that simply using these attributes at test-time does not improve performance, but our training strategy, for example, on the iNaturalist [36] dataset, leads to an average improvement of 4-5% in zero-shot classification accuracy for novel categories of birds [37] and flowers [20]. Similar improvements are observed in domains where a subset of the categories was used to fine-tune the model. By prompting LLMs in various ways, we generate descriptions that capture visual appearance, habitat, and geographic regions and pair them with existing attributes such as the taxonomic structure of the categories. We systematically evaluate their ability to improve zero-shot categorization in natural domains. Our findings suggest that geographic priors can be just as effective and are complementary to visual appearance. Our method also outperforms prior work on prompt-based tuning of VLMs. We plan to release the benchmark, consisting of 7 datasets, which will contribute to future research in zero-shot recognition.
@@ 1 Introduction
On the technical side, we develop methods to train VLMs with “bag-level” supervision. In our dataset a set of images are grouped with a set of descriptions and lack the image-text correspondences. Some of the descriptions may not apply to an image (e.g., the part may be occluded). However, we find that training by stochastically pairing the images and text within a category, followed by a category-level contrastive loss similar to CLIP objective offers robust improvements in performance. Adapting semi-supervised learning approaches such as FixMatch [31] or Knowledge distillation [3] results in minor improvements (§ 5.7). A detailed investigation of the image-text association within a category suggests the model is able to correctly associate the visual attributes with the corresponding text even they are paired randomly during training (Fig. 2).
We systematically evaluate the effectiveness of our method by assessing the zero-shot classification performance on novel classes. We find that simply using these attributes of novel classes generated by LLMs does not improve performance when using CLIP (Tab. 1). However, our training strategy leads to an average improvement of 4-5% in accuracy across 7 datasets, and outperforms baselines (Tab. 2). For natural domains (e.g., iNaturalist and NABirds), we prompt LLMs in various ways to generate descriptions that capture visual appearance, habitat, and geographic regions, and pair them with existing attributes within the dataset, such as taxonomic structure. Our results indicate that geographic priors are equally effective and complementary to visual appearance cues (see Tab. 3). Training on iNaturalist without any bird classes improves the performance of CLIP on CUB by more than 3%, and we observe similar improvements when evaluating across other domains (see Tab. 4). Improvements are consistent across text generated by different LLMs, as well as by humans (see Tab. 5). Our model also results in relative error reduction of 4.1% over CLIP on the challenging NeWT dataset [35].
@@ 2 Related Work
Zero-shot Image Classification using VLMs. Vision Language models (VLMs) [25, 9, 29, 39, 4] learn to associate images with their corresponding text captions. Learning a shared embedding makes them perform exceptionally well in zero-shot classification tasks when paired with an appropriate text such as the class name during test time. FLAVA [29] learns using paired as well as unpaired images and texts using different losses for multimodal and unimodal understanding. ALIGN [9] uses a large number of noisy image-text data by obtaining alt-texts for images and trains using a contrastive loss. CLIP [25] is trained on a smaller and cleaner dataset of image-text pairs using a similar objective function. It employs a vision model and a language model to learn joint embeddings of images and text. While training, it maximizes the similarity between related image-text pairs and minimizes similarity between unrelated pairs. At test time the similarity over captions such “a photo of a [class name]” over all classes in the domain for each image is found. The image is classified to the class with the caption with highest probability. The original paper shows that manual prompt tuning can boost zero-shot classification accuracy.
Generating Better Prompts. Prior work on prompt tuning [11, 45, 28, 5, 10] has focused on improving the text descriptions of classes. For example, CoOp [44] appends learnable context vectors to the class name texts to improve classification. CoCoOp [43] and related methods have also explored prompting the vision encoder simultaneously. While prompt tuning has proven useful for adapting models to a set of “base categories”, its performance on novel categories still falls short of the CLIP baseline.
Another line of research aims at querying LLMs to generate prompts or attributes of categories. CHiLS [21] refines classes based on GPT descriptions (e.g., taxonomic structure) and maps the image to one of the subcategories to improve classification. We also explore the ability to learn taxonmy-based attributes in our work. Menon et al. [18] and CuPL [24] append class-specific attributes obtained from GPT [2] to simple prompts, e.g. “a photo of a [class]” to improve zero-shot performance at test-time, similar to our approach. However, we find that CLIP struggles to recognize nuanced attributes in fine-grained domains where we observe little to no improvement in classification performance, motivating the need for fine-tuning VLMs.
@@
The work most similar to ours is GIST [13], which also utilizes GPT to generate category-specific texts for fine-grained domains. Their method pairs each image with the $n$ most similar texts within the category based on CLIP similarity, and then consolidates them into a caption using GPT. In contrast, our approach stochastically pairs images with text, a strategy we found to be more robust. Notably, our experiments showed that biasing sampling towards similar image and text pairs led to inferior results (see Section 5.7). However, a key difference is their work aims at improving visual features for few-shot learning while we focus on improving both visual and text features for zero-shot learning on novel classes. We also explore training with larger datasets such as iNaturalist and NABirds.
@@
Figure 3: Fine-tuning VLMs to improve zero-shot performance. a) Our framework for \⃝raisebox{-0.9pt}{1} generating fine-grained attributes per class using LLMs, \⃝raisebox{-0.9pt}{2} category-level fine-tuning of VLMs and \⃝raisebox{-0.9pt}{3} evaluating on a series of challenging unseen scenarios. b) We show examples of texts produced in step \⃝raisebox{-0.9pt}{1}.
@@ 3.1 Dataset Generation
What characteristics can be used to differentiate [class] from other [domain] based on just a photo? Provide an exhaustive list of all attributes that can be used to identify the [domain] uniquely. Texts should be of the form “[domain] with [characteristic]”.
Here [class] is the class name for the $K$ classes in the dataset [domain]. Each domain is associated to different datasets, for example for the CUB200 dataset, [domain] is “bird” and for iNaturalist dataset it is “organism”. Appending the domain [domain] is helpful because it provides context about the set of the other classes to distinguish from, and reduces confusion across similar class names in other domains. The LLM produces $l_{k}$ descriptions for category $k$. We append “a photo of [class]” to the generated texts resulting in descriptions of the form “ a photo of a [class] [domain] with [characteristic]” for each class (more details in the experiment section). This results in a set of descriptions $\mathbf{Y^{k}}$ for each category.
We also separately query about the habitat and geographic location of occurrence for the classes in CUB200, Flowers102, NABirds and iNaturalist datasets. For this purpose we use the prompt
@@
We add the texts obtained to the category-level corpus $\mathbf{Y^{k}}$. The impact of these location specific texts and the improvements offered are described in the results section. Examples of visual and habitat descriptions are in the Appendix. We determine the correctness of the texts produced for the CUB dataset by fact-checking texts of a subset of categories with the help of Wikipedia and experts. We also compare the performance using human descriptions when available. However, a challenge is that this manual vetting process does not scale to large datasets such as iNaturalist, and we therefore rely on empirical results to support the utility of the generated text.
@@ 3.3 Evaluation for Zero-shot Classification
CLIPFT 50.89 $\pm$ 0.05 69.75 $\pm$ 0.08 31.15 $\pm$ 0.02 73.68 $\pm$ 0.00 88.32 $\pm$ 0.00
CLIPFT + A 53.36 $\pm$ 0.08 71.76 $\pm$ 0.06 36.47 $\pm$ 0.02 77.05 $\pm$ 0.00 93.75 $\pm$ 0.01
Table 1: Comparison with CLIP ViT-B/32 on zero-shot performance on fine-grained domains. We compare our method CLIPFT + A to the baselines defined in § 4.3. We significantly improve over baseline CLIP evaluated with both “a photo of a [class] [domain]” and LLM attributes. We also fine-tune CLIP with only “a photo of a [class] [domain]” text and compare with our method to show that our improvements are not due to seeing domain-specific images but also by learning correlations between images and fine-grained attributes.
@@
Table 5: Evaluating model trained using GPT4 with texts obtained from other models. Our model consistently improves over pre-trained CLIP when evaluated with texts obtained from different LLM models (GPT3.5 and LLaMA2-7B) as well as GT aggregated captions and “a photo of a [class] [domain]”.
@@ 4.1 Datasets
We use a variety of fine-grained classification datasets including CUB200 (CUB) [37] which has 200 classes, Flowers 102 [20] with 102 classes, Stanford Cars [12] with 196 classes, FGVC Aircrafts [16] which has 100 classes and Food101 [1] with 101 classes. We also apply our method on some coarser datasets including EuroSAT [8] with 10 classes, and ImageNet [27] containing 1000 classes. For all these datasets, we use the first half of the classes (ordered by ids of the original dataset) for training and the second half for zero-shot testing.
@@
NeWT [35] provides a benchmark for a set of 164 complex binary classification tasks in the natural world that extend beyond species classification. These tasks include determining 1) appearance 2) behavior, 3) context, 4) counting and 5) gestalt. NeWT contains 36k images with 200-400 images per task. We randomly select 50 of the 164 tasks to evaluate our trained model. We manually associate two texts for each task, positive and negative. For example, “a photo of a raptor bird which is not on a utility pole” and “a photo of a raptor bird which is on a utility pole”. We show improvements over CLIP (§ 5.5). All details of texts used as well as categories selected are in the Appendix.
@@ 4.2 Implementation Details
For generating category-level texts for training, we utilize the “gpt-4-0613” API. We set the temperature parameter as 0 so that texts generated are deterministic.
For all queries concerning the classes of iNaturalist (both visual and location) we also append the type of organism as well as it’s scientific name in the question. For example, for the class “Bay Laurel” the query for location information is
@@
While fine-tuning using the texts obtained from a LLM, we train for only 15 epochs on each dataset. On iNaturalist we train for only 5 epochs. We use the official train images of our sampled testing classes as our validation set. We find the best model across the checkpoints of both the normal and momentum encoder by finding the accuracy on this set.
@@ 4.3 Baselines
CLIP refers to pre-trained CLIP tested with “a photo of a [class] [domain]” texts like the original paper.
@@
CLIPFT involves fine-tuning CLIP on training classes using “a photo of [class] [domain]” texts and evaluating on test classes using “a photo of [class] [domain]” texts.
@@ 5.1 Comparison with Baselines
In Table 1 we compare our method to three baselines CLIP, CLIP + A and CLIPFT, all evaluated on unseen classes. Here we use the ViT B/32 architecture for all methods. Our method offers considerable improvements over pre-trained CLIP when using “a photo of a [class] [domain]” text and when using GPT generated text. In difficult fine-grained domains such as CUB, Stanford Cars and FGVC Aircrafts pre-trained CLIP does not utilize text attributes generated by GPT resulting in negligible improvement (decrease on Stanford Cars) compared to “a photo of a [class] [domain]” (see CLIP + A vs CLIP). This motivates the need to fine-tune using these attributes, resulting in significant improvement across all datasets. We also compare to fine-tuning CLIP with “a photo of a [class] [domain]” texts (CLIPFT) to show that the improvement our method achieves is not due to just being trained on images of concerned domain.
We compare our method to previous work CLIP-A-self [17] in Table 2. We follow [17] and use the ViT B/16 architecture and only 16 images per class for training. Again we evaluate on unseen classes. Our method outperforms CLIP-A-self significantly across all datasets. Also, our method offers substantial improvement over CLIP, showing that it scales across architectures. We discuss why CLIP-A-self underperforms in detail in the Appendix.
@@ 5.3 Training on External Domains
We now evaluate under more difficult settings. We train and test on different datasets. Note that in every case, we always remove any overlapping classes. For training on iNat and NABirds we use all visual + taxonomy + habitat information. While testing on CUB we use visual + taxonomy + habitat. For Flowers 102 we use visual + habitat information. In Table 4 we show the accuracy on CUB test classes improves considerably when training using NABirds and iNat even though the images of these datasets have a distribution shift w.r.t. CUB. More strikingly, we show that even when we remove all bird classes from iNat we still offer improvement on CUB test classes compared to CLIP + A (52.83 $\rightarrow$ 53.89). Similarly when we remove all plant classes from iNat, we still get improvement on Flowers 102 test set. This proves that our model is also able to generalize well. It is learning to associate fine-grained attributes to images irrespective of the domain differences in training and testing.
@@ 5.4 Using Novel Texts during Test Time
We evaluate how our model would perform in the absence of the LLM used to generate training texts, during test time (Tab 5). We use GPT3.5 turbo (0613) and LLaMA2-7B [33] for generating visual and habitat texts. We show that our method consistently improves performance over pre-trained CLIP for all types of texts explored.  [^1] LLaMA model does not always produce texts in the specified format and thus needs post-processing. The texts formed finally are considerably different grammatically from the sentences our model has been trained on. Please refer to Appendix for examples of texts produced. Our model also improves performance over pre-trained CLIP while using “a photo of a [class] [domain]” texts.
Reed et al. [26] present a dataset of human-labelled captions per image for the CUB dataset. We use these ground truth texts at test time to evaluate our model. The dataset contains 10 captions for every image, which are all visual attributes of the bird in the image. Since many attributes in the captions are repeated for each image as well as across images of the same class, and to limit the size of the text corpus, we randomly select one caption per image of a given class and aggregate them to form a category corpus. In Table 5 last row, we notice that CLIP does better using these image-level GT texts compared to using GPT4 Visual texts which were category-level (row 2). However, our method still outperforms showing that it is able learn meaningful attributes through noisy labels.

### v2

@@ Abstract
The zero-shot performance of existing vision-language models (VLMs) such as CLIP [29] is limited by the availability of large-scale, aligned image and text datasets in specific domains. In this work, we leverage two complementary sources of information—descriptions of categories generated by large language models (LLMs) and abundant, fine-grained image classification datasets—to improve the zero-shot classification performance of VLMs across fine-grained domains. On the technical side, we develop methods to train VLMs with this “bag-level" image-text supervision. We find that simply using these attributes at test-time does not improve performance, but our training strategy, for example, on the iNaturalist [41] dataset, leads to an average improvement of 4-5% in zero-shot classification accuracy for novel categories of birds [42] and flowers [23]. Similar improvements are observed in domains where a subset of the categories was used to fine-tune the model. By prompting LLMs in various ways, we generate descriptions that capture visual appearance, habitat, and geographic regions and pair them with existing attributes such as the taxonomic structure of the categories. We systematically evaluate their ability to improve zero-shot categorization in natural domains. Our findings suggest that geographic priors can be just as effective and are complementary to visual appearance. Our method also outperforms prior work on prompt-based tuning of VLMs. We release the benchmark, consisting of 14 datasets at https://github.com/cvl-umass/AdaptCLIPZS, which will contribute to future research in zero-shot recognition.
@@ 1 Introduction
On the technical side, we develop methods to train VLMs with “bag-level" supervision. In our dataset a set of images are grouped with a set of descriptions and lack the image-text correspondences. Some of the descriptions may not apply to an image (e.g., the part may be occluded). However, we find that training by stochastically pairing the images and text within a category, followed by a category-level contrastive loss similar to CLIP objective offers robust improvements in performance. Adapting semi-supervised learning approaches such as FixMatch [35] or Knowledge distillation [3] results in minor improvements (§ 5.7). A detailed investigation of the image-text association within a category suggests the model is able to correctly associate the visual attributes with the corresponding text even they are paired randomly during training (Fig. 2).
We systematically evaluate the effectiveness of our method by assessing the zero-shot classification performance on novel classes. We find that simply using these attributes of novel classes generated by LLMs does not improve performance when using CLIP (Tab. 1). However, our training strategy leads to an average improvement of 4-5% in accuracy across 12 datasets, and outperforms baselines (Tab. 2). For natural domains (e.g., iNaturalist and NABirds), we prompt LLMs in various ways to generate descriptions that capture visual appearance, habitat, and geographic regions, and pair them with existing attributes within the dataset, such as taxonomic structure. Our results indicate that geographic priors are equally effective and complementary to visual appearance cues (see Tab. 3). Training on iNaturalist without any bird classes improves the performance of CLIP on CUB by more than 3%, and we observe similar improvements when evaluating across other domains (see Tab. 4). Improvements are consistent across text generated by different LLMs, as well as by humans (see Tab. 5). Our model also results in relative error reduction of 4.1% over CLIP on the challenging NeWT dataset [40].
@@ 2 Related Work
Zero-shot Image Classification using VLMs. Vision Language models (VLMs) [29, 11, 33, 45, 6] learn to associate images with their corresponding text captions. Learning a shared embedding makes them perform exceptionally well in zero-shot classification tasks when paired with an appropriate text such as the class name during test time. FLAVA [33] learns using paired as well as unpaired images and texts using different losses for multimodal and unimodal understanding. ALIGN [11] uses a large number of noisy image-text data by obtaining alt-texts for images and trains using a contrastive loss. CLIP [29] is trained on a smaller and cleaner dataset of image-text pairs using a similar objective function. It employs a vision model and a language model to learn joint embeddings of images and text. While training, it maximizes the similarity between related image-text pairs and minimizes similarity between unrelated pairs. At test time the similarity over captions such “a photo of a [class name]" over all classes in the domain for each image is found. The image is classified to the class with the caption with highest probability. The original paper shows that manual prompt tuning can boost zero-shot classification accuracy.
Generating Better Prompts. Prior work on prompt tuning [13, 51, 32, 7, 12] has focused on improving the text descriptions of classes. For example, CoOp [50] appends learnable context vectors to the class name texts to improve classification. CoCoOp [49] and related methods have also explored prompting the vision encoder simultaneously. While prompt tuning has proven useful for adapting models to a set of “base categories", its performance on novel categories still falls short of the CLIP baseline.
Another line of research aims at querying LLMs to generate prompts or attributes of categories. CHiLS [24] refines classes based on GPT descriptions (e.g., taxonomic structure) and maps the image to one of the subcategories to improve classification. We also explore the ability to learn taxonomy-based attributes in our work. Menon et al. [20] and CuPL [28] append class-specific attributes obtained from GPT [2] to simple prompts, e.g. “a photo of a [class]" to improve zero-shot performance at test-time, similar to our approach. However, we find that CLIP struggles to recognize nuanced attributes in fine-grained domains where we observe little to no improvement in classification performance, motivating the need for fine-tuning VLMs.
@@
Two works similar to ours, GIST [15] and I2MVFormer [22], also utilize GPT to generate category-specific texts for fine-grained domains. GIST pairs each image with the $n$ most similar texts within the category based on CLIP similarity and consolidates them into a caption. In contrast, our approach stochastically pairs images with text, a strategy we found to be more robust. Notably, our experiments showed that biasing sampling towards similar image and text pairs led to inferior results (see § 5.7). I2MVFormer uses a LLM to generate class descriptions based on texts provided by annotators and Wikipedia documents. The key differences between their work and ours lie in the use of human effort in their training data generation and starting from scratch resulting in much lower performance compared to ours. Finally, in contrast to both these works, we explore training using visual, habitat, and location information, as well as training on larger datasets such as NABirds and iNaturalist.
@@
[…]
@@ 3.1 Dataset Generation
What characteristics can be used to differentiate [class] from other [domain] based on just a photo? Provide an exhaustive list of all attributes that can be used to identify the [domain] uniquely. Texts should be of the form “[domain] with [characteristic]".
Here [class] is the class name for the $K$ classes in the dataset [domain]. Each domain is associated to different datasets, for example for the CUB200 dataset, [domain] is “bird" and for iNaturalist dataset it is “organism". Appending the domain [domain] is helpful because it provides context about the set of the other classes to distinguish from, and reduces confusion across similar class names in other domains. The LLM produces $l_{k}$ descriptions for category $k$. We append “a photo of [class]" to the generated texts resulting in descriptions of the form “ a photo of a [class] [domain] with [characteristic]" for each class (more details in the experiment section). This results in a set of descriptions $\mathbf{Y^{k}}$ for each category.
We also separately query about the habitat and geographic location of occurrence for the classes in CUB200, Flowers102, NABirds and iNaturalist datasets. For this purpose we use the prompt:
@@
[…]
@@ 3.3 Evaluation for Zero-shot Classification
CLIPFT 50.81 $\pm$ 0.04 69.61 $\pm$ 0.07 31.10 $\pm$ 0.02 73.68 $\pm$ 0.00 88.32 $\pm$ 0.00
CLIPFT + A 53.34 $\pm$ 0.08 71.63 $\pm$ 0.06 36.41 $\pm$ 0.02 77.05 $\pm$ 0.00 93.71 $\pm$ 0.01
Table 1: Comparison with CLIP ViT-B/32 on zero-shot performance on fine-grained domains. We compare our method CLIPFT + A to the baselines defined in § 4.3. We significantly improve over baseline CLIP evaluated with both “a photo of a [class] [domain]" and LLM attributes. We also fine-tune CLIP with only “a photo of a [class] [domain]" text and compare with our method to show that our improvements are not due to seeing domain-specific images but also by learning correlations between images and fine-grained attributes.
@@
Table 5: Evaluating model trained using GPT4 with texts obtained from other models. Our model consistently improves over pre-trained CLIP when evaluated with texts obtained from different LLM models (GPT3.5 and LLaMA2-7B) as well as GT aggregated captions and “a photo of a [class] [domain]".
@@ 4.1 Datasets
We use a variety of fine-grained classification datasets including CUB [42] (200 classes), Flowers 102 [23] (102 classes), Stanford Cars [14] (196 classes), FGVC Aircrafts [18] (100 classes) and Food101 [1] (101 classes). We also apply our method on some coarser datasets including EuroSAT [10] (10 classes), ImageNet [31] (1000 classes), CalTech101 [5] (100 classes), DTD [4] (47 classes), Oxford Pets [26] (37 classes), Sun397 [44] (397 classes) and UCF101 [36] (101 classes). For all these datasets, we use the first half of the classes (ordered by ids of the original dataset) for training and second half for zero-shot testing.
@@
NeWT [40] provides a benchmark for a set of 164 complex binary classification tasks in the natural world that extend beyond species classification. These tasks include determining 1) appearance 2) behavior, 3) context, 4) counting and 5) gestalt. NeWT contains 36k images with 200-400 images per task. We randomly select 50 of the 164 tasks to evaluate our trained model. We manually associate two texts for each task, positive and negative. For example, “a photo of a raptor bird which is not on a utility pole" and “a photo of a raptor bird which is on a utility pole". We show improvements over CLIP (§ 5.5). All details of texts used as well as categories selected are in the Appendix.
@@ 4.2 Implementation Details
For generating category-level texts for training, we utilize the “gpt-4-0613" API. We set the temperature parameter as 0 so that texts generated are deterministic.
For all queries concerning the classes of iNaturalist (both visual and location) we also append the type of organism as well as it’s scientific name in the question. For example, for the class “Bay Laurel" the query for location information is
@@
While fine-tuning using the texts obtained from an LLM, we train for only 15 epochs on each dataset. On iNaturalist, we train for only 5 epochs. We find hyperparameters by splitting the train classes into two equal parts, training on the first half, and validating on the second. We then fix the best hyperparameters found and train on all train classes.
@@ 4.3 Baselines
CLIP refers to pre-trained CLIP tested with “a photo of a [class] [domain]" texts like the original paper.
@@
CLIPFT involves fine-tuning CLIP on training classes using “a photo of [class] [domain]" texts and evaluating on test classes using “a photo of [class] [domain]" texts.
@@ 5.1 Comparison with Baselines
Tab. 1 compares our method to three baselines CLIP, CLIP + A and CLIPFT, all evaluated on unseen classes. Here we use the ViT B/32 architecture for all methods. Our method offers considerable improvements over pre-trained CLIP when using “a photo of a [class] [domain]" text and when using GPT generated text. In difficult fine-grained domains such as CUB, Stanford Cars and FGVC Aircrafts pre-trained CLIP does not utilize text attributes generated by GPT resulting in negligible improvement (decrease on Stanford Cars) compared to “a photo of a [class] [domain]" (see CLIP + A vs CLIP). This motivates the need to fine-tune using these attributes, resulting in significant improvement across all datasets. We also compare to fine-tuning CLIP with “a photo of a [class] [domain]" texts (CLIPFT) to show that the improvement our method achieves is not due to just being trained on images of concerned domain.
We compare our method to previous work CLIP-A-self [19] in Tab. 2. We follow [19] and use the ViT B/16 architecture and only 16 images per class for training. Again we evaluate on unseen classes. Our method outperforms CLIP-A-self significantly across all datasets. Also, our method offers substantial improvement over CLIP, showing that it scales across architectures. We discuss why CLIP-A-self underperforms in detail in the Appendix. We also show results on the 14 datasets benchmark in Appendix.
@@ 5.3 Training on External Domains
We now evaluate under more difficult settings. We train and test on different datasets, always removing any overlapping classes. For training on iNat and NABirds we use all visual + taxonomy + habitat information. While testing on CUB we use visual + taxonomy + habitat. For Flowers 102 we use visual + habitat information. Tab. 4 we show the accuracy on CUB test classes improves considerably when training using NABirds and iNat even though the images of these datasets have a distribution shift w.r.t. CUB. More strikingly, we show that even when we remove all bird classes from iNat we still offer improvement on CUB test classes compared to CLIP + A (52.83 $\rightarrow$ 53.89). Similarly when we remove all plant classes from iNat, we still get improvement on Flowers 102 test set. This proves that our model is also able to generalize well. It is learning to associate fine-grained attributes to images irrespective of the domain differences in training and testing.
@@ 5.4 Using Novel Texts during Test Time
We evaluate how our model would perform in the absence of the LLM used to generate training texts, during test time (Tab 5). We use GPT3.5 turbo (0613) and LLaMA2-7B [38] for generating visual and habitat texts. We show that our method consistently improves performance over pre-trained CLIP for all types of texts explored.  [^1] LLaMA model does not always produce texts in the specified format and thus needs post-processing. The texts formed finally are considerably different grammatically from the sentences our model has been trained on. Please refer to Appendix for examples of texts produced. Our model also improves performance over pre-trained CLIP while using “a photo of a [class] [domain]" texts.
Reed et al. [30] present a dataset of human-labelled captions per image for the CUB dataset. We use these ground truth texts at test time to evaluate our model. The dataset contains 10 captions for every image, which are all visual attributes of the bird in the image. Since many attributes in the captions are repeated for each image as well as across images of the same class, and to limit the size of the text corpus, we randomly select one caption per image of a given class and aggregate them to form a category corpus. In Tab. 5 last row, we notice that CLIP does better using these image-level GT texts compared to using GPT4 Visual texts which were category-level (row 2). However, our method still outperforms showing that it is able learn meaningful attributes through noisy labels.